OpenStudy (anonymous):

sin^2 65 + sin^2 25 + cos^2 35 + cos^2 55 = ?

OpenStudy (vishweshshrimali5):

Use this: 1. \[\large{\sin(90 - \theta) = \cos{\theta}}\] 2. \[\large{\sin^2 x + \cos^2 x = 1}\]

OpenStudy (anonymous):

got it.. thank u :)

OpenStudy (vishweshshrimali5):

Your welcome :)